Hi all, im having trouble with the following problem:
The energy stored in a spring is given by the formula: E = 100x^2 where x is the compression. The rate of change of energy with respect to the compression is called the thrust T of the spring.
How do I:
a) Express the thrust Ta of the spring in terms of x.
b) Find the thrust at x:1.5
Any help would be fantastic.
-Nick

Answer by Alan3354(69443)

The energy stored in a spring is given by the formula: E = 100x^2 where x is the compression. The rate of change of energy with respect to the compression is called the thrust T of the spring.
How do I:
a) Express the thrust Ta of the spring in terms of x.
T = 1st derivative dE/dx = 200x
-------------------------------
b) Find the thrust at x:1.5
= 200*1.5 = 300 (no units given)